OpenStudy (anonymous):

how do i find r in a geometric sequence? 1,1/3,1/9,1/27

OpenStudy (anonymous):

Divide 2nd term / 1st term Divide again to verify 3rd term / 2nd term

OpenStudy (anonymous):

ok well, first off, this isnt a geometric sequence right?

OpenStudy (anonymous):

since its not a common ratio

OpenStudy (anonymous):

3rd/2nd=2nd/1st=r so to know it is geometric we must to devide for arithmetic we have to minus 2nd terms -1st terms=n

OpenStudy (anonymous):

It is a GM There is a common ratio

OpenStudy (anonymous):

well i know that the 2nd term that when multiplied by 3 gives me the 3rd and when done once more gives me the third term, is that enough to say its a geometric sequence, its clear i dont know how to to do 1/3-1

OpenStudy (anonymous):

so its 3?

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 }\div1 = \frac{ 1 }{ 3}\]

OpenStudy (anonymous):

no its 1/3 2nd and 3rd terms 1/9*3/1=1/3

OpenStudy (anonymous):

\[\frac{ 1 }{9 }\div \frac{ 1 }{ 3}=\frac{ 1 }{3 }\]

OpenStudy (anonymous):

-_-im dumb, thnx :)